### v1

@@ 5.4 $4/5$-MMS with Copies
In this section, we present an algorithm which, given an ordered instance, outputs an $\alpha$-MMS allocation with at most $\lfloor n/3\rfloor$ distinct copies for any $\alpha\leq 4/5$.
Similar to Section 5.3, first we reduce the instance but this time by Algorithm 6 using $R^{\alpha}_{0},R^{\alpha}_{1},\ldots,$ and $T^{\alpha}$. Analogously, it is not difficult to see that if $k$ agents are removed (i.e. are assigned a bundle) by the end of the algorithm, then in total at most $\lfloor k/3\rfloor$ duplicated goods are introduced and each good is duplicated at most once. Also, since we remove these $k$ agents and all the goods in these $k$ bags, in the future no more copies of them can be used. Since all these rules are $\alpha$-valid reductions, it suffices to give an $\alpha$-MMS allocation with at most $\lfloor(n^{*}-k)/3\rfloor$ distinct copies for the instance obtained by the remaining agents and goods where $n^{*}$ is the original number of agents. Let $n=n^{*}-k$.
@@
For an arbitrary ordered instance $\mathcal{I}$ and $\alpha\leq 1$, let $\mathcal{I}^{\prime}=\mathtt{T-reduce(\mathcal{I},\alpha)}$. Then for all agents $i\in N^{\prime}$, $\mu_{i}(\mathcal{I}^{\prime})\geq\mu_{i}(\mathcal{I})$.
@@
For the rest of this (sub)-section, let $\alpha\leq 4/5$, $\mathcal{I}$ be an $R^{\alpha}$-irreducible and $T^{\alpha}$-irreducible instance, and $\mu_{i}$ be $\mu_{i}(\mathcal{I})$. Let $n$ be the number of agents right before the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, =$\mathcal{B}=\{B_{1},\ldots,B_{a-1}\}$ be the set of bags allocated during this phase, and $n^{\prime}=n-a+1$ be the number of remaining agents after this phase.
@@
Towards contradiction, assume for some $j>2$, $|B_{j}|\leq 2$ and $B_{j}$ is allocated to agent $i$. Note that since $\mathcal{I}$ is $R_{0}^{\alpha}$-irreducible, $v_{a}(1)<\alpha\cdot\mu_{a}$ for all agents $a$ and hence $|B|\geq 2$ for all $B\in\mathcal{B}$. So for all $\ell\in[n]$, we have $\{\ell,n+\ell\}\in B_{\ell}$ and $B_{j}=\{j,n+j\}$. Note that for all $j^{\prime}<j$, $v_{i}(\{j^{\prime},n+j^{\prime}\})>v_{i}(\{j,n+j\})\geq\alpha\cdot\mu_{i}$. Hence, for all $j^{\prime}<j$, $|B_{j^{\prime}}|=2$. Let $a_{1},a_{2},a_{3}$ be the three agents who received $B_{1},B_{2},B_{3}$ respectively. Note that the precondition of Rule $T^{\alpha}$ is met for these agents which is a contradiction with $\mathcal{I}$ being $T^{\alpha}$-irreducible. ∎
Observation 2.
For $n\geq 2$, $\{B_{1},B_{2}\}\subseteq\mathcal{B}$ and $v_{i}(B_{j})\leq 3\alpha/2\cdot\mu_{i}$ for all $i\in N$ and $j\in[2]$.
@@
Since the instance is $R^{\alpha}_{0}$-irreducible, $v_{i}(2)\leq v_{i}(1)<\alpha\cdot\mu_{i}$. Therefore, $|B_{j}|\geq 2$. Let $g$ be the last good added to $B_{j}$. Since $g\geq n+1$, $v_{i}(g)\leq v_{i}(n+1)<\alpha/2\cdot\mu_{i}$. We have
$v_{i}(B_{j})$ $=v_{i}(B_{j}\setminus\{j\})+v_{i}(B_{j})<\alpha\cdot\mu_{i}+\alpha/2\cdot\mu_{i}=3\alpha/2\cdot\mu_{i}.$
Now assume $B_{2}\notin\mathcal{B}$. Note that for all $B\notin\mathcal{B}$, $v_{i}(B)<\alpha\cdot\mu_{i}$. Otherwise, $B$ should have been allocated during the round robin phase. Therefore
$n\cdot\mu_{i}$ $\leq v_{i}(M)$
@@
$<3\alpha/2\cdot\mu_{i}+(n-1)\alpha\cdot\mu_{i}$
$\leq(4/5n+2/5)\cdot\mu_{i},$ ($\alpha\leq 4/5$)
@@
Let $i$ be a remaining agent after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ (if any). Then |$|\{B\in\mathcal{B}\mid v_{i}(B)>\alpha\cdot\mu_{i}\}|\geq 2n/3$ or $n\leq 5$.
@@
Assume $n>5$. Let +=$\mathcal{B}^{+}=\{B\in\mathcal{B}\mid v_{i}(B)>\alpha\cdot\mu_{i}\}$. By definition, for all $B\in\mathcal{B}\setminus\mathcal{B}^{+}$, $v_{i}(B)\leq\alpha\cdot\mu_{i}$. Let $\bar{\mathcal{B}}$ be the set of $n-|\mathcal{B}|=n^{\prime}$ many remaining bags after the round robin phase. For all $B\in\bar{\mathcal{B}}$, $v_{i}(B)<\alpha\cdot\mu_{i}$. Otherwise, $B$ should have been allocated to $i$ (or some other agent who values it at least $\alpha$ fraction of their MMS value). We consider two cases:
Case 1: $v_{i}(B_{2})\leq 4\alpha/3\cdot\mu_{i}$.
For all $B\in\mathcal{B}\setminus\{B_{1},B_{2}\}$, Lemma 19 implies $|B|\geq 3$ and Lemma 13 implies $v_{i}(B)\leq 4\alpha/3\cdot\mu_{i}$. Also by Observation 2, $v_{i}(B_{1})\leq 3\alpha/2\cdot\mu_{i}$ and we have $v_{i}(B_{2})\leq 4\alpha/3\cdot\mu_{i}$. Therefore
$\sum_{B\in\mathcal{B}^{+}}v_{i}(B)$ $\leq 4\alpha/3\cdot\mu_{i}(|\mathcal{B}^{+}|-1)+3\alpha/2\cdot\mu_{i}$
$\leq\mu_{i}(16|\mathcal{B}^{+}|/15+2/15).$ ($\alpha\leq 4/5$)
@@
$\leq\mu_{i}(16|\mathcal{B}^{+}|/15+2/15)+(|\mathcal{B}|-|\mathcal{B}^{+}|)\alpha\cdot\mu_{i}+(n-|\mathcal{B}|)\alpha\cdot\mu_{i}$
$\leq\mu_{i}(2/15+4n/5+4|\mathcal{B}^{+}|/15).$ ($\alpha\leq 4/5$)
@@
Therefore, $|\mathcal{B}^{+}|\geq 3(n-2)/4\geq 2n/3$.
Case 2: $v_{i}(B_{2})>4\alpha/3\cdot\mu_{i}$.
If $|B_{2}|\geq 3$, then by Lemma 13, $v_{i}(B_{2})\leq 4\alpha/3\cdot\mu_{i}$. Also if $|B_{2}|=1$, $v_{i}(B_{2})<\alpha\cdot\mu_{i}$ since the instance is $R^{\alpha}_{0}$ irreducible. Therefore, $|B_{2}|=2$. It implies $|B_{1}|=2$ and we have $B_{1}=\{1,n+1\}$ and $B_{2}=\{2,n+2\}$. Let $a$ and $b$ be the agents who received $B_{1}$ and $B_{2}$. If $v_{i}(\{1,n+3\})\geq\alpha\cdot\mu_{i}$, then the instance is not $T^{\alpha}$-irreducible since $a,b,i$ satisfy its precondition. Therefore, $v_{i}(1)+v_{i}(n+3)<\alpha\cdot\mu_{i}$.
$v_{i}(1)+v_{i}(n+2)$ $\geq v_{i}(2)+v_{i}(n+2)=v_{i}(B_{2})>4\alpha/3\cdot\mu_{i}.$
Therefore, $v_{i}(n+3)<v_{i}(n+2)-\alpha/3\cdot\mu_{i}<(\alpha/2-\alpha/3)\cdot\mu_{i}=\alpha/6\cdot\mu_{i}$. Now fix an index $j>2$ and let $g$ be the last good added to $B_{j}$. We have
@@
$<\alpha\cdot\mu_{i}+\alpha/6\cdot\mu_{i}$
$<\mu_{i}.$ ($\alpha\leq 4/5$)
Therefore, +=$\mathcal{B}^{+}=\{B_{1},B_{2}\}$ and we have
$n\cdot\mu_{i}$ $\leq v_{i}(M)$
$\leq 3\alpha\cdot\mu_{i}+(|\mathcal{B}|-2)\alpha\cdot\mu_{i}+(n-|\mathcal{B}|)\alpha\cdot\mu_{i}$
$\leq\mu_{i}(4/5+4n/5).$ ($\alpha\leq 4/5$)
We obtain that $n\leq 4$ which is a contradiction with $n>5$. ∎
@@
If $n\geq 6$, for all the remaining agents $i$ after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, and all the remaining bags $B_{j}$, we have $v_{i}(B_{j}\cup\{n^{\prime}\})\geq\alpha\cdot\mu_{i}$.
@@
By Lemma 14, the bags $B_{1},\ldots,B_{n-n^{\prime}}$ are allocated during the round robin phase. By Lemma 20, there exists at least $2n/3\geq n^{\prime}$ many bags $B_{i_{1}},\ldots,B_{i_{n^{\prime}}}$ for $i_{1}<i_{2}<\ldots<i_{n^{\prime}}\leq n-n^{\prime}$ that were allocated to other agents during the round robin phase and $v_{i}(B_{i_{\ell}})>\alpha\cdot\mu_{i}$ for all $\ell\in[n^{\prime}]$. Note that $i_{n^{\prime}}\geq n^{\prime}$. Since $B_{j}$ have been assigned a good in all but maybe the last round, we have $|B_{j}|\geq|B_{i_{n^{\prime}}}|-1$. Let $B_{j}=\{g_{1},\ldots,g_{k}\}$ and $B_{i_{n^{\prime}}}=\{g^{\prime}_{1},\ldots,g^{\prime}_{k^{\prime}}\}$ for $j=g_{1}<\ldots<g_{k}$ and $i_{n^{\prime}}=g^{\prime}_{1}<\ldots<g^{\prime}_{k^{\prime}}$. For all $\ell\in[k^{\prime}-1]$, since $g_{\ell}$ is allocated before $g^{\prime}_{\ell+1}$, $g_{\ell}<g^{\prime}_{\ell+1}$ and hence $v_{i}(g_{\ell})\geq v_{i}(g^{\prime}_{\ell+1})$. We have
@@
returns an $\alpha$-MMS allocation with at most $\lfloor n^{*}/3\rfloor$ many distinct copies.
@@
Let $n$ be the number of agents after applying $\mathtt{T-reduce(\mathcal{I},\alpha)}$. Since $T^{\alpha}$ and $R^{\alpha}_{k}$ are valid reductions for all $k<|M|/n^{*}$, all agents who receive a bag during the $\mathtt{T-reduce(\mathcal{I},\alpha)}$ receive at least $\alpha$ fraction of their MMS value and the MMS value of the remaining agents do not decrease after this process. All the agents who receive a bag during the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}$ get at least $\alpha$ fraction of their value. Let agent $i$ be an agent who receives bag $B_{j}$ after the round robin phase. Let $g$ be the good that was duplicated and added to $B_{j}$ right before allocating it to agent $i$. Since $g\leq n^{\prime}$, $v_{i}(g)\geq v_{i}(n^{\prime})$. By Lemma 21, $v_{i}(B_{j}\cup\{g\})\geq\alpha\cdot\mu_{i}$. Therefore, the final allocation is $\alpha$-MMS.
Now we prove at most $\lfloor n^{*}/3\rfloor$ distinct copies are used. Let $k=n^{*}-n$ be the number of agents who receive a bag in $\mathtt{T-reduce(\mathcal{I},\alpha)}$. Only when applying $T^{\alpha}$, new copies are introduced and one copy per three agents that are removed. Furthermore, the goods that are allocated during $T^{\alpha}$ along with their copies are removed from the instance. Therefore, they cannot be copied more than once. So during $\mathtt{T-reduce(\mathcal{I},\alpha)}$, at most $\lfloor k/3\rfloor$ distinct copies are used. In the duplication phase of $\mathsf{BagFill\mbox{-}RoundRobin}$, we duplicate items $1,2,\ldots,n^{\prime}$ once. By Corollary 6, $n^{\prime}\leq\lfloor n/3\rfloor$. Therefore, in total at most $\lfloor n^{*}/3\rfloor$ distinct copies are used. ∎

### v2

@@ 5.5 $4/5$-MMS with Copies
In this section, we present an algorithm which, given an ordered instance, outputs a simple $\alpha$-MMS allocation with at most $\lfloor n/3\rfloor$ distinct copies for any $\alpha\leq 4/5$.
Similar to Section 5.4, first we reduce the instance but this time by running Algorithm 6 which uses $R^{\alpha}_{0},R^{\alpha}_{1},\ldots,$ and $T^{\alpha}$. Analogous to Algorithm 5, it is not difficult to see that if $k$ agents are removed (i.e. are assigned a bundle) by the end of Algorithm 6, then (i) at most $\lfloor k/3\rfloor$ duplicated goods are introduced, (ii) each good is duplicated at most once, and (iii) all the duplicated goods and their corresponding original instances are allocated to distinct agents. Also, since we remove these $k$ agents and all the goods in these $k$ bags, in the future no more copies of these goods can be used and no other goods can be allocated to these agents. Since all these rules are $\alpha$-valid reductions, it suffices to give a simple $\alpha$-MMS allocation with at most $\lfloor(n^{*}-k)/3\rfloor$ distinct copies for the instance obtained by the remaining agents and goods where $n^{*}$ is the original number of agents. Let $n=n^{*}-k$.
@@
For an arbitrary ordered instance $\mathcal{I}$ and $\alpha\leq 4/5$, let $\mathcal{I}^{\prime}=\mathtt{T-reduce(\mathcal{I},\alpha)}$. Then for all agents $i\in N^{\prime}$, $\mu_{i}(\mathcal{I}^{\prime})\geq\mu_{i}(\mathcal{I})$.
@@
For the rest of this (sub)-section, let $\alpha\leq 4/5$, $\mathcal{I}$ be an $R^{\alpha}$-irreducible and $T^{\alpha}$-irreducible instance, and $\mu_{i}(\mathcal{I})=1$. Let $n$ be the number of agents right before the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, =$\mathcal{B}=\{B_{1},\ldots,B_{a-1}\}$ be the set of bags allocated during this phase, and $n^{\prime}=n-a+1$ be the number of remaining agents after this phase.
@@
Towards contradiction, assume for some $j>2$, $|B_{j}|\leq 2$ and $B_{j}$ is allocated to agent $i$. Note that since $\mathcal{I}$ is $R_{0}^{\alpha}$-irreducible, $v_{a}(1)<\alpha$ for all agents $a$ and hence $|B|\geq 2$ for all $B\in\mathcal{B}$. So for all $\ell\in[n]$, we have $\{\ell,n+\ell\}\in B_{\ell}$ and $B_{j}=\{j,n+j\}$. Note that for all $j^{\prime}<j$, $v_{i}(\{j^{\prime},n+j^{\prime}\})>v_{i}(\{j,n+j\})\geq\alpha$. Hence, for all $j^{\prime}<j$, $|B_{j^{\prime}}|=2$. Let $a_{1},a_{2},a_{3}$ be the three agents who received $B_{1},B_{2},B_{3}$ respectively. Note that the precondition of Rule $T^{\alpha}$ is met for these agents which is a contradiction with $\mathcal{I}$ being $T^{\alpha}$-irreducible. ∎
Observation 3.
For $n\geq 2$, $\{B_{1},B_{2}\}\subseteq\mathcal{B}$ and $v_{i}(B_{j})\leq 3\alpha/2$ for all $i\in N$ and $j\in[2]$.
@@
Since the instance is $R^{\alpha}_{0}$-irreducible, $v_{i}(2)\leq v_{i}(1)<\alpha$ for all $i\in N$. Therefore, $|B_{j}|\geq 2$. Let $g$ be the last good added to $B_{j}$. Since $g\geq n+1$, $v_{i}(g)\leq v_{i}(n+1)<\alpha/2$. We have
$v_{i}(B_{j})$ $=v_{i}(B_{j}\setminus\{g\})+v_{i}(g)<\alpha+\alpha/2=3\alpha/2.$
Now assume $B_{2}\notin\mathcal{B}$. Let $i$ be an agent who is not allocated any bundle in $\mathcal{B}$. Note that for all $B\notin\mathcal{B}$, $v_{i}(B)<\alpha$. Otherwise, $B$ should have been allocated during the round robin phase. Therefore
$n$ $\leq v_{i}(M)$
@@
$<3\alpha/2+(n-1)\alpha$
$\leq(4/5n+2/5)\cdot,$ ($\alpha\leq 4/5$)
@@
Let $i$ be a remaining agent after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ (if any). Then |$|\{B\in\mathcal{B}\mid v_{i}(B)>\alpha\}|\geq 2n/3$ or $n\leq 5$.
@@
Assume $n\geq 6$. Let +=$\mathcal{B}^{+}=\{B\in\mathcal{B}\mid v_{i}(B)>\alpha\}$. By definition, for all $B\in\mathcal{B}\setminus\mathcal{B}^{+}$, $v_{i}(B)\leq\alpha$. Let $\bar{\mathcal{B}}$ be the set of $n-|\mathcal{B}|=n^{\prime}$ many remaining bags after the round robin phase. For all $B\in\bar{\mathcal{B}}$, $v_{i}(B)<\alpha$. Otherwise, $B$ should have been allocated to $i$ (or some other agent who values it at least $\alpha$ fraction of their MMS value). We consider two cases:
Case 1: $v_{i}(B_{2})\leq 4\alpha/3$.
For all $B\in\mathcal{B}\setminus\{B_{1},B_{2}\}$, Lemma 20 implies $|B|\geq 3$ and Lemma 14 implies $v_{i}(B)\leq 4\alpha/3$. Also by Observation 3, $v_{i}(B_{1})\leq 3\alpha/2$ and we have $v_{i}(B_{2})\leq 4\alpha/3$. Therefore
$\sum_{B\in\mathcal{B}^{+}}v_{i}(B)$ $\leq 4\alpha/3(|\mathcal{B}^{+}|-1)+3\alpha/2$
$\leq(16|\mathcal{B}^{+}|/15+2/15).$ ($\alpha\leq 4/5$)
@@
$\leq(16|\mathcal{B}^{+}|/15+2/15)+(|\mathcal{B}|-|\mathcal{B}^{+}|)\alpha+(n-|\mathcal{B}|)\alpha$
$\leq(2/15+4n/5+4|\mathcal{B}^{+}|/15).$ ($\alpha\leq 4/5$)
@@
Therefore, $|\mathcal{B}^{+}|\geq(3n-2)/4\geq 2n/3$ (for $n\geq 6$).
Case 2: $v_{i}(B_{2})>4\alpha/3$.
If $|B_{2}|\geq 3$, then by Lemma 14, $v_{i}(B_{2})\leq 4\alpha/3$. Also if $|B_{2}|=1$, $v_{i}(B_{2})<\alpha$ since the instance is $R^{\alpha}_{0}$ irreducible. Therefore, $|B_{2}|=2$. It implies $|B_{1}|=2$ and we have $B_{1}=\{1,n+1\}$ and $B_{2}=\{2,n+2\}$. Let $a$ and $b$ be the agents who received $B_{1}$ and $B_{2}$. If $v_{i}(\{1,n+3\})\geq\alpha$, then the instance is not $T^{\alpha}$-irreducible since $a,b,i$ satisfy its precondition. Therefore, $v_{i}(1)+v_{i}(n+3)<\alpha$.
$v_{i}(1)+v_{i}(n+2)$ $\geq v_{i}(2)+v_{i}(n+2)=v_{i}(B_{2})>4\alpha/3.$
Therefore,
$v_{i}(n+3)$ $<v_{i}(n+2)-\alpha/3$
$<(\alpha/2-\alpha/3)$ ($v_{i}(n+2)\leq v_{i}(n+1)<\alpha/2$ due to Observation 1)
$=\alpha/6.$
Now fix an index $j>2$ and let $g$ be the last good added to $B_{j}$. We have
@@
$<\alpha+\alpha/6$
$=7\alpha/6.$
Therefore,
$n$ $\leq v_{i}(M)$
$=v_{i}(B_{1})+v_{i}(B_{2})+\sum_{B\in\mathcal{B}\setminus\{B_{1},B_{2}\}}v_{i}(B)+\sum_{B\in\bar{\mathcal{B}}}v_{i}(B)$
$<3\alpha+(|\mathcal{B}|-2)7\alpha/6+(n-|\mathcal{B}|)\alpha$
$\leq(4n/5+(2|\mathcal{B}|+8)/15)$ ($\alpha\leq 4/5$)
$\leq(14n/15+4/15).$ ($|\mathcal{B}|\leq n-2$)
We obtain that $n<4$ which is a contradiction with $n\geq 6$. ∎
@@
If $n\geq 6$, for all the remaining agents $i$ after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, and all the remaining bags $B_{j}$, we have $v_{i}(B_{j}\cup\{n^{\prime}\})\geq\alpha$.
@@
By Lemma 15, the bags $B_{1},\ldots,B_{n-n^{\prime}}$ are allocated during the round robin phase. By Lemma 21, there exists at least $2n/3\geq n^{\prime}$ many bags $B_{i_{1}},\ldots,B_{i_{n^{\prime}}}$ for $i_{1}<i_{2}<\ldots<i_{n^{\prime}}\leq n-n^{\prime}$ that were allocated to other agents during the round robin phase and $v_{i}(B_{i_{\ell}})>\alpha$ for all $\ell\in[n^{\prime}]$. Note that $i_{n^{\prime}}\geq n^{\prime}$. Since $B_{j}$ have been assigned a good in all but maybe the last round, we have $|B_{j}|\geq|B_{i_{n^{\prime}}}|-1$. Let $B_{j}=\{g_{1},\ldots,g_{k}\}$ and $B_{i_{n^{\prime}}}=\{g^{\prime}_{1},\ldots,g^{\prime}_{k^{\prime}}\}$ for $j=g_{1}<\ldots<g_{k}$ and $i_{n^{\prime}}=g^{\prime}_{1}<\ldots<g^{\prime}_{k^{\prime}}$. For all $\ell\in[k^{\prime}-1]$, since $g_{\ell}$ is allocated before $g^{\prime}_{\ell+1}$, $g_{\ell}<g^{\prime}_{\ell+1}$ and hence $v_{i}(g_{\ell})\geq v_{i}(g^{\prime}_{\ell+1})$. We have
@@
returns a simple $\alpha$-MMS allocation with at most $\lfloor n^{*}/3\rfloor$ many distinct copies.
@@
Let $n$ be the number of agents after applying $\mathtt{T-reduce(\mathcal{I},\alpha)}$. Since $T^{\alpha}$ and $R^{\alpha}_{k}$ are valid reductions for all $k<|M|/n^{*}$, all agents who receive a bag during the $\mathtt{T-reduce(\mathcal{I},\alpha)}$ receive at least $\alpha$ fraction of their MMS value and the MMS value of the remaining agents do not decrease after this process. Without loss of generality, let the MMS value of the remaining agents be $1$ at this point of the algorithm. All the agents who receive a bag during the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}$ get at least $\alpha$ fraction of their value. Let agent $i$ be an agent who receives bag $B_{j}$ after the round robin phase. Let $g$ be the good that was duplicated and added to $B_{j}$ right before allocating it to agent $i$. Since $g\leq n^{\prime}$, $v_{i}(g)\geq v_{i}(n^{\prime})$. By Lemma 22, $v_{i}(B_{j}\cup\{g\})\geq\alpha$. Therefore, the final allocation is $\alpha$-MMS.
We now give an upper bound on the total number of copies and prove that the generated MMS allocation is simple, namely that there are $t\leq\lfloor n^{*}/3\rfloor$ distinct copies, and that the $2t$ instances, composed of these copies and their corresponding original goods, are allocated to $2t$ distinct agents. Let $k=n^{*}-n$ be the number of agents who receive a bag in $\mathtt{T-reduce(\mathcal{I},\alpha)}$. Only when applying $T^{\alpha}$, new duplications are introduced and one duplication per three agents that are removed. So during $\mathtt{T-reduce(\mathcal{I},\alpha)}$, at most $\lfloor k/3\rfloor$ distinct copies are used. Each instance of these good is allocated to a different agent. Furthermore, all the instances of the goods that are allocated during $T^{\alpha}$ along with their owners, are removed from the instance. Therefore, these goods cannot be copied more than once and their owners cannot get any other goods later on. In the duplication phase of $\mathsf{BagFill\mbox{-}RoundRobin}$, we duplicate items $1,2,\ldots,n^{\prime}$ once and allocate the original and the duplications to $2n^{\prime}$ different agents. By Corollary 6, $n^{\prime}\leq\lfloor n/3\rfloor$. Therefore, in total at most $\lfloor n^{*}/3\rfloor$ distinct copies are used and the allocation is simple. ∎